OpenStudy (anonymous):

How to solve y^2-2y-8=7 ?

OpenStudy (anonymous):

first start by subtracting 7 from both sides

OpenStudy (anonymous):

How would I go after getting 0=y^2-2y-8-7 ? I was absent when learning this

OpenStudy (anonymous):

you should then get a binomial equal to 0 from here, you can a) factor b) complete the square c)use the quadratic equation ( i would avoid using this unless absolutely necessary)

OpenStudy (anonymous):

well -8-7 = ?

OpenStudy (anonymous):

-15?

OpenStudy (anonymous):

ok so what you now have is 0=y^2-2y-15

OpenStudy (anonymous):

how would i factor from there?

OpenStudy (anonymous):

(x+a)(x+b)=x^2+(a+b)x+ab so start by looking at factors of -15 i suggest you list them

OpenStudy (anonymous):

basically a*b= -15 a+b=-2 in this particular case solve for a and b

OpenStudy (anonymous):

1,-15,15, -1, 3, -5, -3,-5

OpenStudy (anonymous):

ok now which multiple pair, when added together will give you -2?

OpenStudy (anonymous):

Oh I see that now! Its -5 and 3